OpenStudy (anonymous):

differentiate 2x^3+(x^2)(y)-xy^3=2

OpenStudy (baldymcgee6):

Do you know how to start?

OpenStudy (anonymous):

i thought so but it came out to the wrong answer

OpenStudy (baldymcgee6):

Well.. let's see what you did and maybe we can see if you made any mistakes.

OpenStudy (anonymous):

dy/dx (2x^3+x^2y-xy^3)=dy/dx2 2*3x^2 (x^2dy/dx +y)- (xdy/dxy^3+y)=0

OpenStudy (baldymcgee6):